OpenStudy (anonymous):

HELP I can't figure out how to eliminate the parameters. x = 3 sin^3t y = 3 cos^3t

OpenStudy (asnaseer):

are you aware of the identity: \(\cos^2(\theta)+\sin^2(\theta)=1\) ?

OpenStudy (anonymous):

Yes, but I have no idea how to get to that.

OpenStudy (asnaseer):

try squaring each of your expressions and then adding them :)

OpenStudy (anonymous):

I'm sorry, but what?

OpenStudy (dumbcow):

@asnaseer , the sin/cos are cubed functions

OpenStudy (asnaseer):

sorry - I may have misread your question, I read it as \(x=3\sin(3t)\)

OpenStudy (anonymous):

No, it is read as \[3 \sin^3 t \]

OpenStudy (asnaseer):

so are your equations:\[x=3\sin^3(t)\]\[y=3\cos^3(t)\]

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

:)

OpenStudy (dumbcow):

\[\sin t = (\frac{x}{3})^{1/3}\] \[y = 3 \cos t (1-\sin^2 t)\] make the substitution

OpenStudy (anonymous):

I have little to no idea as to how to make it squared and not cubed.

OpenStudy (dumbcow):

sorry \[\cos t = (1-\sin^2 t)^{1/2}\]

OpenStudy (asnaseer):

you can rearrange your equations to become:\[\sin(t)=\sqrt[3]{x/3}\]\[\cos(t)=\sqrt[3]{y/3}\]and then use the cos/sin identity I gave above

OpenStudy (anonymous):

Shouldn't Sine and Cosine be squared to work in the Pythagorean theorem identity?

OpenStudy (asnaseer):

Use either the response given by @dumbcow or use the fact that:\[\cos^2(t)+\sin^2(t)=1\]

OpenStudy (asnaseer):

they are both equivalent

OpenStudy (anonymous):

I know this is redundant, but Sine and Cosine aren't supposed to be squared?

OpenStudy (asnaseer):

do you agree that for any angle \(\theta\):\[\cos^2(\theta)+\sin^2(\theta)=1\]

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

So just let \(\theta=t\) and we get:\[\cos^2(t)+\sin^2(t)=1\]

OpenStudy (asnaseer):

Now all you need to do is to substitute the expressions for \(\cos(t)\) and \(\sin(t)\) that we derived above and you are done :)